# Can Medical Vision-Language Pre-training Succeed with Purely Synthetic Data?

### v1

@@ 2 Related Work
Medical Vision Language Pre-training. Recent work on MedVLP has focused on integrating visual and textual modalities, particularly for chest X-ray (CXR) images. Studies such as (Zhang et al., 2020; Huang et al., 2021; Wang et al., 2022; Liu et al., 2023b; d; c; Wan et al., 2024) have concentrated on aligning CXR images with paired radiology reports. Some methods also leverage external datasets to boost performance, raising concerns about generalizability (Wu et al., 2023; Zhang et al., 2023; Li et al., 2024; Phan et al., 2024a). However, all current MedVLP approaches rely heavily on large-scale, real image-text paired datasets like MIMIC-CXR (Johnson et al., 2019b). Some even require additional human-annotated datasets or manual interventions to improve model performance (Wu et al., 2023; Zhang et al., 2023; Phan et al., 2024a), which limits their scalability and accessibility.
3 Methods
@@ 3.1 Exploring Imperfections in Real Data
For MedVLP, the most commonly used dataset is MIMIC-CXR (Johnson et al., 2019a; b), a collection of chest x-ray (CXR) images paired with their corresponding textual reports. after following the preprocessing steps outlined in previous works (Zhang et al., 2023; Wang et al., 2022; Huang et al., 2021), this dataset provides a total of 213,384 image-text pairs for pre-training. And all images must be frontal views according to the preprocessing steps outlined in (Huang et al., 2021).
Previous work on VLP with natural images (Xu et al., 2023b) has shown that data quality, including image fidelity and long-tailed distribution, significantly impacts model performance. However, the quality of MedVLP datasets remains underexplored due to ambiguity in defining medical image quality, stemming from diverse imaging protocols. Additionally, quantifying data distribution is complex, as radiology reports often describe patterns across multiple anatomical regions rather than distinct categories.To address these challenges, we develop a systematic pipeline to thoroughly analyze the data issues in the MIMIC-CXR (Johnson et al., 2019b) dataset, as detailed in the following sections.
Low-Quality and Mismatched Image-Text Pairs. Our aim is to explore and identify issues related to image quality in the MIMIC-CXR dataset (Johnson et al., 2019a), rather than to completely clean the dataset, as creating a perfect dataset and filtering out all low-quality samples is infeasible for large-scale multimodal datasets (Xu et al., 2023a). Inspired by (Bannur et al., 2023), which highlights various issues with poor-quality images, we design six queries for a Multimodal Large Language Model (MLLM), utilizing the InternVL2-26B model [^2] https://huggingface.co/OpenGVLab/InternVL2-26B (Chen et al., 2023; 2024b). Each CXR image from the MIMIC-CXR dataset is paired with these six queries, and the MLLM process each query independently. The process is depicted in Fig 2 (b).
@@
Uncovering Long-tailed Distribution in MIMIC-CXR. As demonstrated in previous work on natural image-text data (Xu et al., 2023a; Hammoud et al., 2024), a long-tailed distribution in VLP datasets negatively impacts model performance. Therefore, we aim to explore the data distribution of the MIMIC-CXR dataset. However, directly evaluating the text distribution at the sample level, as done in (Xu et al., 2023a), is challenging because each radiology report often describes multiple patterns or anatomical regions, unlike natural image captions that typically focus on a single object (Zhang et al., 2024).
Instead, we adopt an alternative approach by using an off-the-shelf Named Entity Recognition (NER) tool to extract all medical entities, treating them as representatives of the report’s concepts and exploring the dataset distribution at the entity level. For this, we use RaTE [^4] https://huggingface.co/Angelakeke/RaTE-NER-Deberta (Zhao et al., 2024), a model specifically designed for NER tasks on radiology reports. RaTE automatically classifies the extracted entities into five categories: [ABNORMALITY, NON-ABNORMALITY, DISEASE, NON-DISEASE, ANATOMY]. We display the top 50 frequent entiites distribution of each entity type in Fig 2 (c). We display the top 50 frequent entiites distribution of each entity type in Fig 7,8,11,9,10. As shown, all entity types exhibit a severe long-tailed distribution. The MIMIC-CXR (Johnson et al., 2019b) includes a total of 154,049 unique entities, with 55,047 Abnormality, 36,365 Non-Abnormality, 23,017 Disease, 22,103 Non-Disease, and 40,517 Anatomy entities.
@@
CXR Report Generation. To generate the synthetic reports, the pipeline is depicted in Fig 6. We select a general LLM, Llama3.1-70B-Instruct [^5] https://huggingface.co/meta-llama/Meta-Llama-3.1-70B-Instruct as the report generator, and we extensively ablate the performance of the report generator with other LLMs in Fig 4. We query the LLM using prompts that include the entity list, as shown in Fig 6.
@@
Here, $f(e)$ denotes the current frequency of entity $e$ in the dataset. This ensures a balanced distribution of entities across the synthetic dataset.
@@ 3.3 Synthetic Data Training for MedVLP
Finally, we use the synthetic dataset, SynCXR, to train a MedVLP model and explore how effectively a model can learn from pure synthetic data. Since there are many existing methods for MedVLP, we select simple baseline models like ConVIRT (Zhang et al., 2020) and GLoRIA (Huang et al., 2021) for the following reasons:
@@
GLoRIA (Huang et al., 2021) extends ConVIRT by incorporating both global and regional contrastive learning to train the encoders on paired medical images and reports.
These models are open-source, straightforward, and minimize the influence of external factors on evaluating synthetic data for MedVLP. For retraining these two methods on our synthetic dataset, SynCXR, we strictly use their official codebases [^7] https://github.com/marshuang80/gloria [^8] https://github.com/edreisMD/ConVIRT-pytorch. More complex models may introduce unnecessary complications.
Excluding Complex Models. Recent models like BioViL (Boecking et al., 2022) and BioViL-T (Bannur et al., 2023) lack publicly available training code, making them impractical for re-training with synthetic data. Knowledge-enhanced MedVLP models such as MedKLIP, KAD, and MAVL (Wu et al., 2023; Zhang et al., 2023; Phan et al., 2024b) rely on external tools and human-annotated data to incorporate additional knowledge, making direct implementation with synthetic data challenging and introducing unnecessary variables.
@@ 4.1 Downstream Task Datasets and Configurations
For downstream tasks, we evaluate the effectiveness of synthetic data for MedVLP across four tasks. Details on the datasets and implementation are provided in Appendix, Sec A.
Zero-shot Medical Image Classification. Following the guidelines in (Phan et al., 2024b; Wu et al., 2023), we perform this task on seven datasets: CheXpert (Saporta et al., 2022), ChestXray-14 (Wang et al., 2017), PadChest-seen, PadChest-unseen, PadChest-rare (Bustos et al., 2020), RSNA (Shih et al., 2019), and SIIM (Steven G. Langer & George Shih, 2019), using the dataset splits from (Phan et al., 2024b). Evaluation metrics include AUC, F1, and ACC.
Zero-shot Medical Image Visual Grounding. In line with (Phan et al., 2024b), this task is conducted on the RSNA (Shih et al., 2019), SIIM (Steven G. Langer & George Shih, 2019), and Covid-19 Rural (Desai et al., 2020) datasets, using official splits and metrics. Grounding performance is evaluated with IoU, and Dice score.
Medical Image Fine-tuned Classification. As described in (Phan et al., 2024b), we use the RSNA (Shih et al., 2019), SIIM (Steven G. Langer & George Shih, 2019), Covid-19 CXR-2 (Pavlova et al., 2022), and ChestXray-14 (Wang et al., 2017) datasets. During fine-tuning, all model parameters, including the pre-trained vision encoder and linear classifier, are updated. The AdamW optimizer is applied with a learning rate of $1\times 10^{-4}$, batch size of 64, and training runs for 50 epochs. Evaluation follows the AUC score protocol in (Huang et al., 2021; Wang et al., 2022; Zhou et al., ).
Medical Image Fine-tuned Segmentation. This task uses the RSNA (Shih et al., 2019), SIIM (Steven G. Langer & George Shih, 2019), and Covid-19 Rural (Desai et al., 2020) datasets, following preprocessing from (Wang et al., 2022; Huang et al., 2021). U-Net (Ronneberger et al., 2015) is used for fine-tuning, freezing the pre-trained vision encoder and updating only the decoder parameters. Performance is measured using the Dice score, adhering to the evaluation protocol from (Huang et al., 2021).
@@ 4.2 Experimental Results
Zero-shot Visual Grounding. We further evaluate the effectiveness of synthetic data in improving MedVLP models’ local visual understanding capabilities through zero-shot grounding tasks. Tab 3(b) presents the performance of zero-shot grounding on RSNA (Shih et al., 2019), Covid-19 Rural (Desai et al., 2020), and SIIM (Steven G. Langer & George Shih, 2019). Across all datasets, MedVLP models pretrained on the SynCXR dataset achieve superior performance compared to those trained on the real dataset, with an average increase of 1.42% IoU and 0.97% Dice scores. The mixed dataset further enhances performance, with 4.06% IoU and 2.92% Dice scores on average. This demonstrates that the SynCXR dataset not only benefits global cross-modal feature learning but also improves local visual understanding for MedVLP models.
@@ 5 Analysis
According to Tab 4(b), for both MedVLP methods, using real images with synthetic reports results in decreased performance, likely due to the persistent long-tailed distribution, as the synthetic reports are generated based on real images. However, using real reports with synthetic images slightly improves performance, as synthetic images can be curated using our image filtering procedure to ensure high quality, avoiding issues commonly found in real datasets. Using both synthetic images and synthetic reports achieves the highest performance, indicating that a well-curated synthetic dataset can significantly enhance MedVLP performance.
@@
Impact of Different Report Generators. We also examine the impact of using different LLMs for synthetic report generation. As shown in Fig 4 (Bottom Left), we compare two general LLMs, LLaMA 3.1 (8B and 70B), and two medical-specific LLMs, Meditron3 (8B [^10] https://huggingface.co/OpenMeditron/Meditron3-8B and 70B [^11] https://huggingface.co/OpenMeditron/Meditron3-70B). Despite Meditron3 being trained specifically on medical corpora and inheriting weights from LLaMA, the dataset generated by LLaMA 3.1-70B-Instruct achieves the best performance. This indicates that a powerful general LLM is effective for generating synthetic datasets, and using domain-specific fine-tuned versions may degrade the quality of the synthetic data.
Impact of Different Image Generators. We evaluate various text-to-image models for synthetic CXR image generation, including CXR-IRGen (Shentu & Al Moubayed, 2024), LLM-CXR (Lee et al., 2023), and RoentGen (Bluethgen et al., 2024). As shown in Fig 4 (Bottom Right), datasets generated by RoentGen lead to the best performance for both MedVLP methods. This is likely because RoentGen is the only image generation model verified by clinicians, suggesting that the quality of image generation models is crucial for building synthetic datasets, and models should be validated by clinical experts.
@@ 6 Conclusion
Appendix A Downstream Tasks Configuration
A.1 Dataset Details
@@
CheXpert (Irvin et al., 2019) contains 224,316 chest X-rays from 65,240 patients at Stanford Hospital, with an official validation set of 200 studies and a test set of 500 studies, both annotated by board-certified radiologists. Our evaluation on the five observations in the official test set follows protocols from earlier studies (Tiu et al., 2022b; Irvin et al., 2019).
SIIM (Steven G. Langer & George Shih, 2019) consists of over 12,000 frontal X-rays annotated with pneumothorax masks, split into training (60%), validation (20%), and test (20%) sets.
@@
A.2 Implementation Details
Zero-shot Image Classification. The CXR images undergo a two-step preprocessing: resizing to $256\times 256$, followed by center cropping to $224\times 224$. As per (Huang et al., 2021), pixel values are normalized to $[0,1]$. The processed image is passed through a visual encoder and projector to generate the image embedding $\hat{\mathbf{v}}_{i}$. Simultaneously, the text prompts are processed through a text encoder to obtain text embeddings $\hat{\mathbf{l}}_{i}$. Classification is based on cosine similarity between image and text embeddings. If the similarity between the image embedding and the positive prompt (e.g., disease) is higher than that with the negative prompt (e.g., No disease), the classification is positive, and vice versa. The prompt design follows (Tiu et al., 2022a) for both ConVIRT and GLoRIA.
@@ A.2 Implementation Details
Medical Image Fine-tuned Segmentation. For segmentation tasks on the RSNA (Shih et al., 2019), SIIM (Steven G. Langer & George Shih, 2019), and Covid-19 Rural (Wang et al., 2020) datasets, we fine-tune both the pre-trained vision encoder and decoder. Training is performed with early stopping at 50 epochs, using a learning rate of 2e-4 and weight decay of 0.05. AdamW is the optimizer, with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. Batch sizes are 8 for SIIM and 16 for RSNA. All configurations follow the protocol from (Huang et al., 2021).
Appendix B Extra Visualization

### v2

@@ Abstract
Can Medical Vision-Language Pre-training Succeed with Purely Synthetic Data?
Che Liu1,2, Zhongwei Wan3, Haozhe Wang4, Yinda Chen5, Talha Qaiser2, Chen Jin2, Fariba Yousefi2, Nikolay Burlutskiy2, Rossella Arcucci1 1Imperial College London, UK, 2AstraZeneca, UK, 3Ohio State University, US, 4Hong Kong University of Science and Technology, 5University of Science and Technology of China che.liu21@imperial.ac.uk
@@ 2.1 Exploring Imperfections in Real Data
For MedVLP, the most commonly used dataset is MIMIC-CXR (Johnson et al., 2019a, b), a collection of chest x-ray (CXR) images paired with their corresponding textual reports. after following the preprocessing steps outlined in previous works (Zhang et al., 2023; Wang et al., 2022; Huang et al., 2021), this dataset provides a total of 213,384 image-text pairs for pre-training. And all images must be frontal views according to the preprocessing steps outlined in (Huang et al., 2021).
Previous work on VLP with natural images (Xu et al., 2023a) has shown that data quality, including image fidelity and long-tailed distribution, significantly impacts model performance. However, the quality of MedVLP datasets remains underexplored due to ambiguity in defining medical image quality, stemming from diverse imaging protocols. Additionally, quantifying data distribution is complex, as radiology reports often describe patterns across multiple anatomical regions rather than distinct categories.To address these challenges, we develop a systematic pipeline to thoroughly analyze the data issues in the MIMIC-CXR (Johnson et al., 2019b) dataset, as detailed in the following sections.
Low-Quality and Mismatched Image-Text Pairs. Our aim is to explore and identify issues related to image quality in the MIMIC-CXR dataset (Johnson et al., 2019a), rather than to completely clean the dataset, as creating a perfect dataset and filtering out all low-quality samples is infeasible for large-scale multimodal datasets (Xu et al., 2023b). Inspired by (Bannur et al., 2023), which highlights various issues with poor-quality images, we design six queries for a Multimodal Large Language Model (MLLM), utilizing the InternVL2-26B model [^2] https://huggingface.co/OpenGVLab/InternVL2-26B (Chen et al., 2023, 2024b). Each CXR image from the MIMIC-CXR dataset is paired with these six queries, and the MLLM process each query independently. The process is depicted in Fig 2 (b). We described the queries for each function in detail in Sec. B.
@@
Uncovering Long-tailed Distribution in MIMIC-CXR. As demonstrated in previous work on natural image-text data (Xu et al., 2023b; Hammoud et al., 2024), a long-tailed distribution in VLP datasets negatively impacts model performance. Therefore, we aim to explore the data distribution of the MIMIC-CXR dataset. However, directly evaluating the text distribution at the sample level, as done in (Xu et al., 2023b), is challenging because each radiology report often describes multiple patterns or anatomical regions, unlike natural image captions that typically focus on a single object (Zhang et al., 2024).
Instead, we adopt an alternative approach by using an off-the-shelf Named Entity Recognition (NER) tool to extract all medical entities, treating them as representatives of the report’s concepts and exploring the dataset distribution at the entity level. For this, we use RaTE [^4] https://huggingface.co/Angelakeke/RaTE-NER-Deberta (Zhao et al., 2024), a model specifically designed for NER tasks on radiology reports. RaTE automatically classifies the extracted entities into five categories: [ABNORMALITY, NON-ABNORMALITY, DISEASE, NON-DISEASE, ANATOMY]. We display the top 50 frequent entiites distribution of each entity type in Fig 2 (c). We display the top 50 frequent entiites distribution of each entity type in Fig 7,8,11,9,10. As shown, all entity types exhibit a severe long-tailed distribution. As shown, all entity types exhibit a severe long-tailed distribution in the MIMIC-CXR (Johnson et al., 2019b), which contains 154,049 unique entities, with 55,047 Abnormality, 36,365 Non-Abnormality, 23,017 Disease, 22,103 Non-Disease, and 40,517 Anatomy entities.
@@
CXR Report Generation. To generate the synthetic reports, the pipeline is depicted in Fig 6. We select a general LLM, Llama3.1-70B-Instruct as the report generator, and we extensively ablate the performance of the report generator with other LLMs in Fig 4. We query the LLM using prompts that include the entity list, as shown in Fig 6.
@@
Here, $f(e)$ denotes the current frequency of entity $e$ in the dataset. This ensures a balanced distribution of entities across the synthetic dataset. We set $k=9$, $m=3$, and $\tau_{\text{max}}=15$ in our work during the generation stage.
@@ 2.3 Synthetic Data Training for MedVLP
In this work, we use the synthetic dataset, SynCXR, to train a MedVLP model and investigate how effectively a model can learn from purely synthetic data. Given the abundance of existing MedVLP methods, we focus on simple baseline models for the following reasons:
@@
GLoRIA (Huang et al., 2021) extends ConVIRT by adding both global and regional contrastive learning, enabling more effective training of encoders on paired medical images and reports.
These models are open-source, straightforward, and minimize the influence of external factors, which is crucial for evaluating synthetic data in the context of MedVLP. For retraining these models on our synthetic dataset, SynCXR, we adhere strictly to their official codebases [^6] https://github.com/marshuang80/gloria [^7] https://github.com/edreisMD/ConVIRT-pytorch. We are aware that recent methods (Boecking et al., 2022; Bannur et al., 2023; Wu et al., 2023; Zhang et al., 2023; Phan et al., 2024b) either lack publicly available code or rely on additional human annotations, which make direct implementation with synthetic data challenging and introduce unwanted variables.
@@ 3.1 Downstream Task Datasets and Configurations
For downstream tasks, we evaluate the effectiveness of synthetic data for MedVLP across four tasks. We strictly follow the downstream setting described in Phan et al. (2024b) to evaluate our method. Due to the page limit, detailed information on the datasets and implementation is provided in Appendix, Sec. C.
@@ 3.2 Experimental Results
Zero-shot Visual Grounding. We further evaluate the effectiveness of synthetic data in improving MedVLP models’ local visual understanding capabilities through zero-shot grounding tasks. Tab 3(b) presents the performance of zero-shot grounding on RSNA (Shih et al., 2019), Covid-19 Rural (Desai et al., 2020), and SIIM (Steven G. Langer and George Shih, 2019). Across all datasets, MedVLP models pretrained on the SynCXR dataset achieve superior performance compared to those trained on the real dataset, with an average increase of 1.42% IoU and 0.97% Dice scores. The mixed dataset further enhances performance, with 4.06% IoU and 2.92% Dice scores on average. This demonstrates that the SynCXR dataset not only benefits global cross-modal feature learning but also improves local visual understanding for MedVLP models.
@@ 4 Analysis
According to Tab 4(b), for both MedVLP methods, using real images with synthetic reports results in decreased performance, likely due to the persistent long-tailed distribution, as the synthetic reports are generated based on real images. However, using real reports with synthetic images slightly improves performance, as synthetic images can be curated using our image filtering procedure to ensure high quality, avoiding issues commonly found in real datasets. Using both synthetic images and synthetic reports achieves the highest performance, indicating that a well-curated synthetic dataset can significantly enhance MedVLP performance. Furthermore, We evaluate the MedVLP method on the cleaned MIMIC-CXR dataset, as detailed in Section D and Table 5
@@
Impact of Different Report Generators. We also examine the impact of using different LLMs for synthetic report generation. As shown in Fig 4 (Bottom Left), we compare two general LLMs, Llama 3.1 (8B and 70B), and two medical-specific LLMs, Meditron3 (8B and 70B) OpenMeditron (2025). Despite Meditron3 being trained specifically on medical corpora and inheriting weights from Llama, the dataset generated by Llama 3.1-70B-Instruct achieves the best performance. This indicates that a powerful general LLM is effective for generating synthetic datasets, and using domain-specific fine-tuned versions may degrade the quality of the synthetic data.
Impact of Different Image Generators. We evaluate various text-to-image models for synthetic CXR image generation, including CXR-IRGen (Shentu and Al Moubayed, 2024), LLM-CXR (Lee et al., 2023), and RoentGen (Bluethgen et al., 2024). As shown in Fig 4 (Bottom Right), datasets generated by RoentGen lead to the best performance for both MedVLP methods. This is likely because RoentGen is the only image generation model verified by clinicians, suggesting that the quality of image generation models is crucial for building synthetic datasets, and models should be validated by clinical experts.
@@ 5 Conclusion
Limitation
While our work demonstrates that MedVLP can successfully operate using purely synthetic data, there are several limitations to consider. Firstly, the success of the synthetic data approach heavily relies on the capabilities of the language and image generation models. Moreover, the synthetic data generation process requires a filtering stage, which introduces additional computational overhead. Although MedVLP is capable of handling noisy data—an issue also present in real-world datasets—the imperfect pairing of synthetic data may still present challenges. Evaluating the realism of synthetic data through human judgment is valuable but costly. In future, we aim to design more efficient data filtering methods and develop metrics that can better simulate human evaluation to enhance data quality.
Appendix A Related Work
Medical Vision Language Pre-training. Recent work on MedVLP has focused on integrating visual and textual modalities, particularly for chest X-ray (CXR) images. Studies such as (Zhang et al., 2020; Huang et al., 2021; Wang et al., 2022; Liu et al., 2023b, d, c; Wan et al., 2024) have concentrated on aligning CXR images with paired radiology reports. Some methods also leverage external datasets to boost performance, raising concerns about generalizability (Wu et al., 2023; Zhang et al., 2023; Li et al., 2024; Phan et al., 2024a). However, all current MedVLP approaches rely heavily on large-scale, real image-text paired datasets like MIMIC-CXR (Johnson et al., 2019b). Some even require additional human-annotated datasets or manual interventions to improve model performance (Wu et al., 2023; Zhang et al., 2023; Phan et al., 2024a), which limits their scalability and accessibility.
Appendix B Queries for Using MLLM to Assess Issued CXR Images
This section provides detailed queries used to guide the MLLM in assessing issued CXR images. Each query is designed to evaluate specific aspects of the images, ensuring their quality and suitability for diagnostic purposes.
Appendix C Downstream Tasks Configuration
C.1 Dataset Details
@@
CheXpert (Irvin et al., 2019) contains 224,316 chest X-rays from 65,240 patients at Stanford Hospital, with an official validation set of 200 studies and a test set of 500 studies, both annotated by board-certified radiologists. Our evaluation on the five observations in the official test set follows protocols from earlier studies (Tiu et al., 2022a; Irvin et al., 2019).
SIIM (Steven G. Langer and George Shih, 2019) consists of over 12,000 frontal X-rays annotated with pneumothorax masks, split into training (60%), validation (20%), and test (20%) sets.
@@
C.2 Implementation Details
Zero-shot Image Classification. The CXR images undergo a two-step preprocessing: resizing to $256\times 256$, followed by center cropping to $224\times 224$. As per (Huang et al., 2021), pixel values are normalized to $[0,1]$. The processed image is passed through a visual encoder and projector to generate the image embedding $\hat{\mathbf{v}}_{i}$. Simultaneously, the text prompts are processed through a text encoder to obtain text embeddings $\hat{\mathbf{l}}_{i}$. Classification is based on cosine similarity between image and text embeddings. If the similarity between the image embedding and the positive prompt (e.g., disease) is higher than that with the negative prompt (e.g., No disease), the classification is positive, and vice versa. The prompt design follows (Tiu et al., 2022b) for both ConVIRT and GLoRIA.
@@ C.2 Implementation Details
Medical Image Fine-tuned Segmentation. For segmentation tasks on the RSNA (Shih et al., 2019), SIIM (Steven G. Langer and George Shih, 2019), and Covid-19 Rural (Wang et al., 2020) datasets, we fine-tune both the pre-trained vision encoder and decoder. Training is performed with early stopping at 50 epochs, using a learning rate of 2e-4 and weight decay of 0.05. AdamW is the optimizer, with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. Batch sizes are 8 for SIIM and 16 for RSNA. All configurations follow the protocol from (Huang et al., 2021).
Appendix D Results on Cleaned MIMIC-CXR
We re-pretrained ConVIRT on the cleaned version of the MIMIC-CXR dataset, and the results are presented in Table 5. As shown, the VLP model pre-trained on the cleaned MIMIC-CXR dataset achieves slightly better performance than the model trained on the original, uncleaned dataset. However, it still falls short when compared to the model pre-trained on our synthetic SynCXR dataset. This performance gap can be attributed to two key factors:
  • Despite filtering out a large number of low-quality samples from the real dataset, it remains challenging to completely remove all poor or unpaired samples. Consequently, some problematic samples persist in the dataset, negatively affecting the VLP process.
  • The cleaning process for MIMIC-CXR primarily targeted image quality, but did not address the long-tailed distribution of entities in the dataset. Simply downsampling or oversampling image-text pairs is not a viable solution. This issue limits the representational balance of the cleaned dataset and impacts overall model performance.
Method Pre-training Data CheXpert ChestXray-14 PadChest-seen RSNA SIIM
AUC $\uparrow$ F1 $\uparrow$ AUC $\uparrow$ F1 $\uparrow$ AUC $\uparrow$ F1 $\uparrow$ AUC $\uparrow$ F1 $\uparrow$ AUC $\uparrow$ F1 $\uparrow$
ConVIRT MIMIC-CXR (cleaned) 53.85 36.45 53.80 13.25 63.51 14.73 80.15 56.10 65.70 44.10
ConVIRT SynCXR 59.49 40.51 56.07 15.43 63.43 15.10 82.08 58.38 75.55 57.43
Table 5: Performance comparison of ConVIRT pre-trained on different datasets.
Appendix E Extra Visualization